OpenStudy (anonymous):

write an equation in point-slope form for the line through the given point with the given slope (5, 2); m = 3 a. y+ 2= 3(x-5) b. y + 2 = 3x-5 c. y- 2 = 3(x+ 5) d. y - 2= 3(x-5)

OpenStudy (anonymous):

@BunnyBree help me? :3

OpenStudy (midhun.madhu1987):

The point-slope of a line is given by: y - y1 = m(x - x1) where m = slope.

OpenStudy (anonymous):

Okay, the point slope formula is: y - y1 = m(x - x1) So, 2 would be y1. 5 would be x1. And of course, m = 3. Now just plug em' in. (:

OpenStudy (anonymous):

ohhh :o

OpenStudy (anonymous):

can you guys help me on some moreee? :3